OpenStudy (anonymous):

Exact value of cos 11pi/12 - using half angle identities

OpenStudy (solomonzelman):

\(\Huge\color{black}{ \cos(\frac{θ}{2})=± \sqrt{\frac{1+ \cos θ}{2} } }\)

OpenStudy (solomonzelman):

\(\Huge\color{black}{ \cos(\frac{330}{2})=± \sqrt{\frac{1+ \cos 330}{2} } }\)